OpenStudy (anonymous):

Two groups of students order burritos and tacos at a local restaurant. One order of 3 burritos and 4 tacos cost $11.33. The other order of 9 burritos and 5 tacos cost $23.56.

OpenStudy (anonymous):

what is the question haha

OpenStudy (radar):

Karnack (Johnny Carson fans) knows. What are the price each for burritos or tacos! Let b = cost of burritos Let t = cost of tacos 3b + 4t = $11.33 given by problem 9b + 5t = $23.56 also given by problem. You have a system of 2 equations with 2 unk. Multiply first equation by 3 getting: 9b + 12t = $33.99 9b + 5t = $23.56 (2nd equation) Now subtract getting: ------------------ 7t =10.43 t (taco) = $1.49 Leaving it at that, for you to substitute the value for t and solve the b (burrito) cost.

OpenStudy (radar):

@Darealmelojr Welcome to open study, come back when more help is needed. Good luck with your studies. Enjoy the Burrito.